# Should we pre-train a decoder in contrastive learning for dense prediction tasks?

### v1

@@ 5 Discussion
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation (Tab. 6). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 5). We showed that pre-training an FPN DeCon adaptation of SlotCon leads to better results when using deep supervision (DeCon-ML) compared to the baselines as well as DeCon-SL, both when transferring encoder only and when transferring encoder and decoder. Framework specific hyperparameter tuning of the auxiliary layers could increase further the performance, as shown in Supplementary material Section 3.
@@
We introduced a combined loss where encoder and decoder loss contribute to the total loss in a weighted manner. We observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a construction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. We also observed in Tab. 3 that dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML).
@@
As shown in Tab. 1, DeCon-SL adaptation of an SSL framework also allows for pre-training a decoder efficiently. However, transferring the FPN decoder along the encoder for VOC object detection did not increase detection performance compared to transferring only the encoder, potentially indicating a form of over-fitting of the FPN decoder to the pre-training dataset. This is alleviated by deep supervision in DeCon-ML adaptation as shown in Tab. 3. We showed that DeCon can be used to adapt two different SSL frameworks using different concepts to learn local information (pixel level VS object/group level). This suggests that the proposed DeCon adaptation can be used in any SSL contrastive method designed for dense downstream tasks.
While DeCon-SL adaptation with an FCN decoder benefits DenseCL framework, the added value of the FPN decoder transfer in DeCon-SL adaptation of DenseCL was not obvious on object detection and instance segmentation downstream tasks. However, the added value of the FPN decoder transfer in DeCon-SL adaptation of SlotCon was present across all scenarios, enhancing model performance, possibly due to the region-based contrastive mechanism of SlotCon. We could not investigate the impact of DeCon-ML adaptation on DenseCL due to its extensive memory requirements, which can be attributed to its reliance on a large number of negative pairs.
DeCon’s performance on VOC object detection and Cityscapes semantic segmentation is constrained by its base SSL framework, SlotCon, preventing it from surpassing the state-of-the-art (PixCon) for these tasks(Tab. 6). As an adaptation, DeCon’s effectiveness largely depends on the underlying framework. While integrating DeCon with the latest state-of-the-art framework could set a new benchmark across all datasets, we were unable to implement PixCon due to the unavailability of its code [37].
DeCon adaptation of an SSL framework was also advantageous for fine-tuning on small scale datasets and out-of-domain dense tasks compared to the encoder-only baseline framework (see Tab. 7). This benefit was also observed when pre-training directly on the target dataset. In such cases, DeCon is valuable compared to an encoder-only pre-training whether we transfer the pre-trained decoder or not.
One limitation of our approach is its high memory and computational demand. Pre-training a decoder using DeCon can significantly increase the parameter count and training time, which can restrict this approach to certain SSL frameworks. For example, the DeCon-SL adaptation of SlotCon with FPN decoder could not fit in a single 80GB memory GPU. More details on the parameter count are available in Supplementary Material Section 4. Another limitation of our work is that we only considered ResNet-50 backbone; however, we expect achieving superior results when using more efficient backbones such as EfficientNet [42].
@@
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed SSL adaptation to offer a multi-stage continual pre-training in a contrastive manner.
@@ 6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that not-only does this adaptation provide a solution for pre-training decoders, but it also improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating these enhancements into an existing framework, we demonstrated that DeCon outperforms or matches state-of-the-art methods across multiple downstream dense prediction tasks.

### v2

@@ 5 Discussion
Table 7: COCO and COCO+ pretraining downstream performance improvement with DeCon-ML-L and Cohen’s d value. PixCon performance was extracted from the original paper.
Method COCO obj det. COCO Inst. Seg. VOC Sem. Seg. City Sem. Seg.
$\Delta$AP $\Delta$AP $\Delta$mIoU $\Delta$mIoU
COCO Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.50 +0.40 -0.40 +0.10
PixCon (2024) (Comp. to SlotCon) 0.00 +0.06 +1.30 +0.51
DeCon-ML-L (Comp. to SlotCon) +0.37 +0.32 +1.42 +0.50
Cohen’s d (DeCon-ML-L VS SlotCon) 2.35 1.94 6.37 2.51
COCO+ Pre-training Performance Improvement $\Delta$ (calculated from the respective papers)
SlotCon (2022) (Comp. to prev. works) +0.60 +0.50 Not avail. Not avail.
PixCon (2024) (Comp. to SlotCon) -0.50 -0.50 -0.20 +0.40
DeCon-ML-L (Comp. to SlotCon) +0.45 +0.27 +1.43 +0.57
Cohen’s d (DeCon-ML-L VS SlotCon) 5.42 3.21 5.25 1.22
@@
In this paper, we introduced a contrastive framework adaptation by incorporating a decoder contrastive loss to an existing framework. We showed that our proposed SSL adaptation outperforms previous state-of-the-art methods on COCO object detection and instance segmentation and achieves comparable results on Pascal VOC semantic segmentation when pre-trained on COCO dataset, and consistently outperforms previous methods when pre-trained on COCO+, and ImageNet except for Cityscapes semantic segmentation (Tab. 1). This adaptation enhances the representation of the pre-trained encoder, improving its transfer performance for dense downstream tasks, regardless of the homogeneity between pre-training and fine-tuning decoders (as shown in Tab. 1).
We introduced a combined loss where encoder and decoder losses contribute to the total loss in a weighted manner.
Figure 2: Pre-training loss dynamics for DeCon-ML-L and SlotCon.
Figure 2 illustrates the loss behavior across different pre-trainings when both the encoder and decoder are used. As shown, decoder pre-training does not influence the encoder’s loss dynamics, supporting the fact that the encoder and decoder loss terms are conceptually non-competing. We also observed that increasing the weight for the decoder loss compared to that of the encoder enhances the performance of the pre-trained models on downstream tasks. This could be attributed to the non-competing nature of an encoder and decoder loss compared to approaches that utilize a reconstruction loss for a decoder and a different loss for the encoder in an encoder-decoder architecture [28]. Additionally, Fig. 3 presents the slots learned by SlotCon encoder and by the different decoder levels of DeCon-ML. It demonstrates that the decoder learns features that capture similar concepts as the encoder’s, while being more spatially precise.
Figure 3: Slots, as defined in SlotCon, learned at different outputs of our architecture. The top left image from COCO validation dataset is overlayed with the slot that was the most represented in original SlotCon’s encoder output feature map, resized to the image’s shape. Other slots displayed are the slots that had the biggest overlap with SlotCon’s main encoder output slot. More details on the slots creation are available in Supplementary Material 6.
@@
Moreover, we noticed that the encoder loss is optional when pre-training with DeCon-ML, demonstrating the efficiency of the decoder deep supervision and channel dropout combination to ensure a comprehensive pre-training of both encoder and decoder. The lack of skip-connections in DeCon-SL with FCN decoder makes the combined encoder-decoder loss preferable to ensure the best pre-training. In Supplementary Material Section 7, we provide experiments on using DeCon-SL with an FPN decoder. We observed that with an FPN decoder, dropout does not provide the same benefits in DeCon-SL compared to DeCon-ML. Hence, channel dropout should only be applied to the different encoder levels when pre-training with a skip-connected decoder and decoder deep supervision (DeCon-ML). Framework specific hyperparameter tuning of the decoder auxiliary layers could also increase the fine-tuning performance (shown in Supplementary Material Section 8).
We demonstrated that our method is effective with a different type of encoders. Specifically, pre-training a ConvNeXt-small backbone, which is both more efficient and larger than the ResNet-50 backbone, also resulted in superior downstream performance when integrated with a decoder in the DeCon-ML-L framework. These results attest to the scalability and practical utility of the proposed approach for pre-training modern backbone architectures.
@@
DeCon was also used to successfully adapt three different SSL frameworks using different concepts (pixel level VS object/group level) to learn local information (Tab. 3). This suggests that the proposed DeCon adaptation could be used in other SSL contrastive methods designed for dense downstream tasks.
Table 8: Pre-training parameter count, GPU setup, and time required per epoch for the different architectures. We study various pre-training setups on the COCO dataset using the SlotCon framework. Module sizes are given in millions of parameters.
Pre-training Enc. size Dec. size Enc. Aux. Layers Size Dec. Aux. Layers Size Full Archi. Size GPU setup Time per epoch (s)
SlotCon 23.51 x 2 0 71.51 0 118.52 1 * A100 80GB 255
DeCon-SL-FCN ($\alpha$ = 0.25) 23.51 x 2 5.31 x 2 71.51 56.83 185.97 1 * A100 80GB 300
DeCon-ML-S ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 59.08 110.03 1 * A100 80GB 280
DeCon-ML-L ($\alpha$ = 0) 23.51 x 2 3.35 x 2 0 227.31 281.02 8 * H100 80GB 165
DeCon adaptation of an SSL framework provided notable advantages for fine-tuning on small-scale datasets and out-of-domain dense prediction tasks, outperforming the encoder-only baseline (see Tab. 4). This improvement persisted even when pre-training was conducted directly on the small target datasets. Furthermore, DeCon achieved superior performance compared to related frameworks when only a limited fraction (5–25%) of the original labeled data was available for fine-tuning. These results highlight the value of DeCon in settings with very limited annotated samples or small-scale, domain-specific datasets, such as those in medical or agricultural applications, regardless of whether the pre-trained decoder is transferred.
Performance of deep learning models can vary across multiple runs due to inherent stochasticity. To mitigate the impact of randomness while considering computational resource constraints, each experiment was conducted three times, with all fine-tunings for a specific downstream task performed on similar hardware. The performance improvements achieved with our DeCon framework compared to the original framework are statistically significant, as evidenced by large Cohen’s $d$ values (>0.8>0.8> 0.8) in Tab. 7, exceeding two standard deviations (see Supplementary Material Section 9), and $p$-values less than 0.05 (0.012 for COCO object detection and 0.048 for instance segmentation) from Wilcoxon signed-rank tests. Details on the computation of these values are provided in Supplementary Material 10. Furthermore, the performance gains shown in Tab. 7 are comparable to those reported in prior studies.
One limitation of our approach is its high memory and computational demand. However, we have shown in Tab. 1 that if GPU is a constraint, the user can decide to reduce the number of decoder levels used in the pre-training and tune the decoder projector layers to reduce the parameter count of DeCon-ML architecture and match the parameter count of the original framework (as shown in Tab. 8), while still improving the performance of the pre-training.
@@
Brempong et al. [4] followed a staged approach, first pre-training the encoder with supervision on annotated datasets, then pre-training the decoder with denoising. However, this multi-stage process relies on labeled data. It could also result in suboptimal feature alignment and increased training complexity. Future works could involve extending our proposed DeCon adaptation to offer a multi-stage continual pre-training in a contrastive manner.
@@ 6 Conclusion
In this paper, we introduced DeCon—a novel contrastive framework adaptation that enhances the performance of self-supervised learning approaches for dense predictions tasks, such as object detection, semantic segmentation, and instance segmentation. We showed that this adaptation improves the encoder’s representation significantly. We proposed two variants of DeCon: DeCon-SL and DeCon-ML. The former introduces a decoder contrastive loss, and the latter extends DeCon-SL by introducing channel dropout and decoder deep supervision to maximize the encoder’s pre-training power. By integrating the proposed contrastive joint encoder-decoder pre-training strategy into an existing framework, we show that DeCon consistently outperforms or matches state-of-the-art methods across diverse dense prediction tasks, highlighting the effectiveness of this unified pre-training approach.
Acknowledgments
Computing resources were provided by the University of British Columbia Cloud Innovation Centre, powered by Amazon Web Services https://cic.ubc.ca/. This research was also partly enabled by support provided by the Digital Research Alliance of Canada https://alliancecan.ca/, and the Research Computing Services group at the University of Calgary https://rcs.ucalgary.ca/RCS_Home_Page.